### v3

# Merlin: Multi-Stage Curriculum Alignment for Multilingual Encoder-LLM Integration in Cross-Lingual Reasoning

Abstract
Large language models (LLMs) excel in English but still struggle with complex reasoning in many low-resource languages (LRLs). Existing methods align LLMs with multilingual encoders, such as LangBridge and MindMerger, raising the accuracy for mid and high-resource languages, yet large performance gap remains for LRLs. We present Merlin, a model-stacking framework that iteratively refines in 2-stages based on a curriculum strategy (from general to specific where general is bilingual bitext and specific is task-specific data) and adapts only a small set of DoRA weights. On the AfriMGSM benchmark Merlin improves exact-match accuracy by +12.9 pp over MindMerger and outperforms GPT-4o-mini by 15.2 pp. It also yields consistent gains on MGSM and MSVAMP (+0.9 and +2.8 pp), demonstrating effectiveness across both low and high-resource settings.
Merlin: Multi-Stage Curriculum Alignment for Multilingual Encoder-LLM Integration in Cross-Lingual Reasoning
Kosei Uemura1   David Guzmán2   Quang Phuoc Nguyen3   Jesujoba Oluwadara Alabi4 En-shiun Annie Lee3,1David Ifeoluwa Adelani2,5 1University of Toronto  2Mila - Quebec AI Institute, McGill University  3Ontario Tech University 4Saarland University  5Canada CIFAR AI Chair {k.uemura,enshiun.lee}@mail.utoronto.ca quangphuoc.nguyen@ontariotechu.net jalabi@lsv.uni-saarland.de {david.guzman, david.adelani}@mila.quebec
1 Introduction
Large language models (LLMs) have achieved impressive results across a broad range of natural language processing (NLP) tasks (openai2024gpt4technicalreport; jiang2024mixtralexperts; geminiteam2024gemini15unlockingmultimodal). Despite these advances, their performance often deteriorates significantly when dealing with low-resource languages (LRLs) (adelani2024irokobenchnewbenchmarkafrican; uemura-etal-2024-afriinstruct; ojo2024goodlargelanguagemodels). This shortfall is partly due to the limited online presence of LRLs both in data availability and NLP tools, and heavy focus on high or mid-resource languages during the pretraining of these models (touvron2023llama2openfoundation; ranathunga_neural_2023; aryabumi2024aya).
To improve reasoning capabilities for LRLs, the most straightforward approach is to adapt LLMs through continual pre-training (alves2024tower; ng2025sea; buzaaba2025lugha) or post-training (ogundepo2025improving). However, these methods often depend on access to in-domain data, which is scarce for LRLs due to the lack of annotators, speakers, and linguistic experts. Additionally, fine-tuning LLMs—especially those with billions of parameters—incurs substantial computational costs and energy usage, making these approaches less feasible in resource-constrained settings (dettmers2023qloraefficientfinetuningquantized; han2024parameterefficientfinetuninglargemodels). LangBridge (yoon-etal-2024-langbridge) and MindMerger (huang2024mindmergerefficientboostingllm) address this challenge by leveraging rich multilingual representations from an external encoder as input to decoder-only LLMs. This design significantly enhances multilingual reasoning capabilities while substantially reducing both data and computational requirements. On the other hand, curriculum learning—a training paradigm in which models are exposed to data of increasing difficulty—shown to improve both convergence speed and final performance (soviany_curriculum_2022) and have been applied in recent works to enhance LLMs reasoning ability (ma-etal-2025-problem).
Figure 1: Overview of Merlin (Multilingual Embedding-Enhanced Reasoning for Language Integration Network), our two-stage framework for multilingual reasoning. In the Mapping Stage, a lightweight mapping layer is trained sequentially on three datasets—starting with general bilingual translation, followed by query-pair translation, and finally task-specific QA translation—to align the multilingual encoder’s outputs with the LLM’s embedding space. In the Embedding Enhancement Stage, this mapping “connector” remains frozen while the LLM body is fine-tuned via parameter-efficient PEFT on QA data, thereby strengthening cross-lingual reasoning.
In this paper, we introduce Multilingual Embedding-Enhanced Reasoning for Language Integration Network (Figure 1), which is abbreviated to Merlin. It is a two-stage framework that integrates insights from model stacking with question alignment. This design combines data-efficient supervised training with targeted adaptation of the LLM’s internal representations, yielding tighter encoder–decoder alignment without the cost of full-model retraining. Inspired by the principles of curriculum learning, in the first stage, we learn a lightweight mapping layer to align a strong multilingual encoder with a frozen LLM through three successive tasks starting with general domain bilingual data, followed by math domain bilingual data, and finally the task data. In the second stage, we fine-tune the LLM with parameter-efficient methods while keeping the mapping layer fixed, enabling the LLM to internalize and better exploit the learned representations during reasoning. We evaluate Merlin on various math reasoning datasets such as Multilingual Grade School Math Benchmark (MGSM) (shi2022languagemodelsmultilingualchainofthought), Multilingual Semantic VAriant Math Problems (MSVAMP) (chen2024breakinglanguagebarriersmultilingual), and AfriMGSM (adelani2024irokobenchnewbenchmarkafrican), an extension of MGSM to several low-resource African languages. Merlin surpasses the strongest baseline, MindMerger, LayAlign, SLAM and QAlign, by more than +11.0 accuracy points on AfriMGSM and delivers consistent gains on MGSM with +2.8 performance gain to MindMerger, demonstrating its effectiveness for cross-lingual mathematical reasoning in multilingual settings, including LRLs. To assess generalization beyond mathematical reasoning, we evaluate Merlin on natural language inference tasks using AfriXNLI (adelani2024irokobenchnewbenchmarkafrican), with +1.6 improvement in performance over MindMerger.
Model Bn Th Sw Ja Zh De Fr Ru Es En MRL HRL Avg
MGSM
Open Models
Metamath 7B 6.4 2.0 4.4 36.4 40.4 52.8 50.4 46.0 58.4 63.6 4.3 49.7 36.1
Gemma 2 9B 56.4 66.4 64.4 65.2 69.6 70.4 59.2 67.6 72.0 76.0 62.4 68.6 66.7
Key Multilingual Approaches–Metamath 7B & mT5‑XL
LangBridge 42.8 50.4 43.2 40.0 45.2 56.4 50.8 52.4 58.0 63.2 45.5 52.3 50.2
QAlign-LoRA 32.4 39.6 40.4 44.0 48.4 54.8 56.8 52.4 59.6 68.0 37.5 54.9 49.6
SLAM 32.0 44.0 40.0 46.0 48.4 54.0 55.2 54.8 56.8 64.8 38.7 54.3 49.6
LayAlign 1.6 3.6 8.0 23.2 23.2 37.2 37.6 34.8 38.4 43.2 4.4 33.9 25.1
MindMerger 43.6 44.4 42.5 45.6 45.6 56.8 58.0 56.0 54.0 64.0 43.5 54.3 51.1
Merlin 43.2 49.2 44.8 48.4 48.8 53.2 55.6 58.0 59.2 63.6 45.7 55.3 52.4
$\Delta$ (Merlin–$M$) -$\mathord{-}0.4$ $\mathord{+}4.8$ $\mathord{+}2.3$ $\mathord{+}2.8$ $\mathord{+}3.2$ -$\mathord{-}3.6$ -$\mathord{-}2.4$ $\mathord{+}2.0$ $\mathord{+}5.2$ -$\mathord{-}0.4$ $\mathbf{\mathord{+}2.2}$ $\mathbf{\mathord{+}1.0}$ $\mathbf{\mathord{+}1.3}$
Key Multilingual Approaches–Gemma 2 9B & NLLB‑600M‑distilled
QAlign-LoRA 58.0 60.8 62.0 56.4 62.4 68.8 68.0 68.0 76.8 80.8 60.3 68.7 66.2
SLAM 63.2 61.2 52.4 62.4 64.4 68.8 53.6 70.0 69.2 69.2 58.9 65.4 63.4
MindMerger 66.8 72.8 69.2 66.8 72.8 78.0 75.2 73.6 76.4 82.4 69.6 75.0 73.4
Merlin 74.8 76.0 77.6 71.6 72.8 73.2 74.8 78.4 77.6 85.6 76.1 76.3 76.2
$\Delta$ (Merlin–$M$) $\mathord{+}8.0$ $\mathord{+}3.2$ $\mathord{+}8.4$ $\mathord{+}4.8$ 0.0 -$\mathord{-}4.8$ -$\mathord{-}0.4$ $\mathord{+}4.8$ $\mathord{+}1.2$ $\mathord{+}3.2$ $\mathbf{\mathord{+}6.5}$ $\mathbf{\mathord{+}1.3}$ $\mathbf{\mathord{+}2.8}$
MSVAMP
Open Models
Metamath 7B 11.8 13.1 11.8 40.9 48.2 56.1 58.0 54.8 59.5 63.0 12.2 54.4 41.7
Gemma 2 9B 63.9 72.0 71.2 78.7 81.8 78.7 79.3 78.3 80.3 80.7 69.0 79.7 76.5
Key Multilingual Approaches–Metamath 7B & mT5‑XL
LangBridge 46.8 46.3 42.1 45.5 50.4 58.1 57.0 55.8 56.9 60.6 45.1 54.9 52.0
QAlign-LoRA 41.7 47.7 54.8 58.0 55.7 62.8 63.2 61.1 63.3 65.3 48.1 61.3 57.4
SLAM 49.1 50.6 55.4 60.6 63.1 64.6 65.4 64.1 66.6 65.3 51.7 64.2 60.5
LayAlign 7.3 15.4 16.9 41.7 45.9 56.4 55.3 52.9 55.0 59.3 13.2 52.4 40.6
MindMerger 51.1 47.9 53.0 54.7 53.6 58.4 60.7 57.7 60.5 64.1 50.7 58.5 56.2
Merlin 53.1 55.6 58.9 62.0 62.3 65.5 65.7 62.2 64.9 68.2 55.9 64.4 61.8
$\Delta$ (Merlin–$M$) $\mathord{+}2.0$ $\mathord{+}7.7$ $\mathord{+}5.9$ $\mathord{+}7.3$ $\mathord{+}8.7$ $\mathord{+}7.1$ $\mathord{+}5.0$ $\mathord{+}4.5$ $\mathord{+}4.4$ $\mathord{+}4.1$ $\mathbf{\mathord{+}5.2}$ $\mathbf{\mathord{+}5.9}$ $\mathbf{\mathord{+}5.6}$
Key Multilingual Approaches–Gemma 2 9B & NLLB‑600M‑distilled
QAlign-LoRA 69.9 72.0 75.5 77.9 77.1 82.3 80.8 80.7 81.8 84.6 72.5 80.7 78.3
SLAM 60.5 70.2 68.8 74.7 74.5 74.5 74.7 73.6 75.7 72.9 66.5 74.4 72.0
MindMerger 69.1 73.1 76.8 79.3 78.3 82.6 79.5 79.5 82.5 82.3 73.0 80.6 78.3
Merlin 72.1 76.1 77.4 78.7 79.8 81.5 81.8 79.8 81.8 83.3 75.2 81.0 79.2
$\Delta$ (Merlin–$M$) $\mathord{+}3.0$ $\mathord{+}3.0$ $\mathord{+}0.6$ -$\mathord{-}0.6$ $\mathord{+}1.5$ -$\mathord{-}1.1$ $\mathord{+}2.3$ $\mathord{+}0.3$ -$\mathord{-}0.7$ $\mathord{+}1.0$ $\mathbf{\mathord{+}2.2}$ $\mathbf{\mathord{+}0.4}$ $\mathbf{\mathord{+}0.9}$
Table 1: Performance on MGSM and MSVAMP. Bn/Th/Sw are mid‑resource (MRL); the other seven languages are high‑resource (HRL). Sub‑headings (e.g. “Metamath‑7B & mT5‑XL”) specify the decoder base model and the multilingual encoder. LangBridge, LayAlign, MindMerger, and Merlin use both components; SLAM and QAlign adopt only the listed decoder. $\Delta$ rows show Merlin’s absolute gain over MindMerger ($\mathord{+}x$ for improvements, -$\mathord{-}x$ for drops).
2 Methodology
Given a frozen multilingual encoder and a lightweight connector into the LLM’s embedding space, Merlin improves cross-lingual reasoning through a two-stage process. The first stage learns to align encoder outputs with the LLM representation space, while the second introduces a small adaptation inside the decoder to internalize this alignment for task solving. The subsections below describe the objectives and data used in each stage.
2.1 Stage I: Model Stacking
Stage I is designed to help the decoder learn to interpret encoder representations without modifying its original parameters. This is accomplished through a three-step curriculum that gradually improves cross-lingual alignment, laying the groundwork for the subsequent task specialization stage, where the model’s reasoning capabilities are further refined.
Problem Definition: Let $q$ be a target-language sentence of length $\ell$. A pretrained multilingual encoder produces $X=\operatorname{Enc}(q)\in\mathbb{R}^{\ell\times d_{\text{enc}}}$. A two-layer perceptron $f_{\sigma}:\mathbb{R}^{d_{\text{enc}}}\!\rightarrow\!\mathbb{R}^{d_{\text{llm}}}$, with parameters $\sigma$, projects these states into the LLM space, $X_{f}=f_{\sigma}(X)$.
The model stacking stage is divided into three consecutive sub-stages that progressively specialize the connector while the encoder and the LLM remain frozen.
(a) General bilingual mapping (Map.)
An initial bridge is established with a generic bitext from $L_{\text{tgt}}$, a generic sentence in the target language, to English (huang2024mindmergerefficientboostingllm). The decoder receives only the projected encoder states
$M_{\sigma}(q)=\bigl[\langle\texttt{bos}\rangle;\,X_{f};\,\langle\texttt{sep}\rangle\bigr],$
and the connector parameters are obtained from
[…]
(b) Question Alignment (zhu2024questiontranslationtrainingbetter) (Align.)
The connector is then refined with translated prompts $(\tilde{q},q^{\star})$, where $\tilde{q}$ is a question in $L_{\text{tgt}}$ and $q^{\star}$ is its English counterpart:
[…]
This step encourages the representations of non-English prompts to align closely with the semantic space utilized by the LLM for English reasoning.
(c) Task-aware Augmentation (Aug.)
Finally, translated question–answer pairs $(\tilde{q},a^{\star})$ inject task-specific supervision. Here the decoder is allowed to see both the encoder projection and its own token embeddings $T(q)=E_{\text{LLM}}[q]$:
$(X_{f},T)(q)=\bigl[\langle\texttt{bos}\rangle;\,X_{f};\,\langle\texttt{sep}\rangle;\,T(q)\bigr].$
The objective becomes
[…]
yielding the final connector $\sigma^{\ast}=\sigma^{(3)}$.
2.2 Stage II: Task Specialization
Stage II internalizes the cross-lingual signal inside the decoder; enhances the integration of multilingual representations within the decoder and plays a critical role in improving performance on complex reasoning tasks. The connector $\sigma^{\ast}$, the encoder and all pretrained LLM weights remain fixed; only rank-constrained DoRA (liu2024doraweightdecomposedlowrankadaptation) matrices $\theta$ inside the decoder are activated. For every English task example $(q,a^{\star})$ the input takes the augmented form $(X_{f},T)(q)$ defined above. The adaptation parameters are obtained through
$\theta^{\ast}=\arg\!\min_{\theta}\mathbb{E}_{(q,a^{\star})}\!\bigl[-\log p_{\text{LLM}_{\theta}}\!\bigl(a^{\star}\mid(X_{f},T)(q)\bigr)\bigr].$
DoRA constrains learning to a negligible fraction of the total parameter count while letting the model gradually absorb how to exploit $X_{f}$ during generation.
Inference.
Given a target-language query $q$, the encoder produces $X$, the connector adds the prefix $X_{f}=f_{\sigma^{\ast}}(X)$, and the adapted decoder $\operatorname{LLM}_{\theta^{\ast}}$ generates the answer in English from $(X_{f},T)(q)$ zero-shot with respect to $L_{\text{tgt}}$.
Model amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul Avg
closed models
GPT-4o-mini 31.6 6.0 56.0 33.6 48.0 25.6 29.2 39.2 44.8 36.8 70.8 15.6 7.6 32.4 45.6 43.6 35.4
GPT-4o 57.6 8.8 64.8 57.6 60.4 51.2 51.6 61.2 58.4 60.8 78.8 31.2 28.0 52.4 62.0 57.2 52.6
open models
MetaMath-Mistral 7B 1.2 2.4 2.0 2.0 2.4 3.6 4.4 2.0 1.2 2.4 10.4 2.8 2.8 1.6 1.6 1.6 2.8
Gemma 2 9B 26.4 4.8 35.2 11.6 22.4 11.6 17.6 9.6 26.0 22.0 61.6 9.2 3.6 20.8 13.6 21.2 19.8
Key multilingual approaches-MetaMath Mistral 7B & AfriTeVa
LangBridge 31.6 2.0 41.2 14.0 28.4 4.0 11.2 17.2 30.0 29.6 54.4 2.8 0.8 18.4 21.2 27.2 20.9
QAlign-LoRA 4.4 3.6 3.2 1.6 5.6 7.6 3.2 4.8 5.6 4.8 18.4 2.4 3.6 4.0 4.4 4.8 5.1
SLAM 1.6 2.0 1.6 1.6 1.2 2.0 2.4 1.6 1.2 4.0 2.4 4.4 2.0 2.8 2.4 1.6 2.2
LayAlign 30.0 6.4 36.4 27.2 32.0 16.0 21.6 26.8 31.6 30.0 42.8 4.0 6.0 23.6 30.4 29.6 24.6
MindMerger 34.4 12.0 54.4 33.6 43.2 22.0 30.0 28.8 40.4 42.8 60.0 8.4 7.2 32.0 41.2 40.8 33.2
Merlin 45.6 12.8 60.8 37.6 48.8 26.4 33.6 39.6 44.0 44.4 58.8 7.6 9.2 38.0 43.2 41.2 37.0
$\Delta$ (Merlin–$M$) $\mathord{+}11.2$ $\mathord{+}0.8$ $\mathord{+}6.4$ $\mathord{+}4.0$ $\mathord{+}5.6$ $\mathord{+}4.4$ $\mathord{+}3.6$ $\mathord{+}10.8$ $\mathord{+}3.6$ $\mathord{+}1.6$ -$\mathord{-}1.2$ -$\mathord{-}0.8$ $\mathord{+}2.0$ $\mathord{+}6.0$ $\mathord{+}2.0$ $\mathord{+}0.4$ $\mathbf{\mathord{+}3.8}$
Key multilingual approaches-Gemma2 9B & NLLB-600M distilled
QAlign-LoRA 45.2 9.6 44.8 29.6 30.8 19.2 21.6 14.8 31.2 30.4 65.6 12.4 6.4 26.8 26.0 30.4 33.7
SLAM 51.6 26.0 43.2 33.6 41.6 37.2 33.6 45.6 40.0 33.2 57.6 21.2 18.8 30.4 41.2 34.8 39.0
MindMerger 45.2 32.4 44.8 34.4 43.2 38.4 36.4 40.0 44.8 36.8 53.2 24.4 21.2 33.6 34.8 40.4 37.7
Merlin 59.2 41.2 60.8 52.0 52.0 49.6 46.8 32.0 58.8 56.0 76.8 37.6 26.0 51.2 54.0 55.6 50.6
$\Delta$ (Merlin–$M$) $\mathord{+}14.0$ $\mathord{+}8.8$ $\mathord{+}16.0$ $\mathord{+}17.6$ $\mathord{+}8.8$ $\mathord{+}11.2$ $\mathord{+}10.4$ -$\mathord{-}8.0$ $\mathord{+}14.0$ $\mathord{+}19.2$ $\mathord{+}23.6$ $\mathord{+}13.2$ $\mathord{+}4.8$ $\mathord{+}17.6$ $\mathord{+}19.2$ $\mathord{+}15.2$ $\mathbf{\mathord{+}12.9}$
Table 2: AfriMGSM performance across various LLMs and key multilingual approaches. We report Exact Match scores averaged over the displayed languages. We use the NLLB-600M distilled model as the encoder, and Gemma2-9b-it for LayAlign, MindMerger, Merlin, and as the base model for QAlign. We use AfriTeVa-v2-Large as the encoder of LangBridge. Note that the released LangBridge source code does not accept NLLB architecture.
Model eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul Avg
closed models
GPT-4o-mini 86.2 80.3 52.2 37.2 64.8 57.2 56.8 31.5 51.7 61.3 63.0 56.2 63.7 47.5 43.2 64.5 54.5 62.0 54.2
GPT-4o 89.2 82.3 71.8 45.0 75.2 68.2 68.0 32.7 69.8 71.2 71.3 71.8 71.5 55.8 52.7 72.0 64.5 67.5 64.3
open models
Aya-101 67.0 59.7 64.2 43.2 57.0 55.5 54.3 33.5 51.7 51.5 55.7 52.2 56.5 47.0 36.7 55.2 54.5 55.3 51.5
Gemma 2 9B 55.3 50.7 43.2 35.3 47.0 40.7 40.2 32.8 38.8 37.8 42.3 40.2 46.3 37.2 35.2 42.5 41.3 44.5 40.3
Key Multilingual Approaches-Gemma2 & NLLB 600m Distilled
QAlign-LoRA 81.0 74.7 51.3 35.7 55.8 49.6 43.3 34.3 41.5 39.7 48.5 48 60.7 43.8 35.0 51.5 47.5 49.3 49.5
SLAM 67.0 61.8 48.2 42.7 47.2 51.3 41.7 32.2 47.5 47.7 46.5 47.2 52.8 38.8 35.7 49.8 52.2 48.3 47.7
MindMerger 80.8 76.3 69.2 60.5 70.8 68.5 66.8 34.5 70.3 54.7 75.2 71.2 71.3 63.7 59.5 73.8 68.8 74.3 67.1
Merlin 90.8 86.7 74.8 60.0 73.3 71.7 68.8 34.8 69.2 59.3 71.3 72.0 72.2 65.5 62.2 73.7 69.2 70.2 68.7
Table 3: AfriXNLI performance on various LLMs, MindMerger, and Merlin. Metric is Accuracy. Average computed on only African languages. NLLB 600M distilled is used as the encoder.
3 Experimental Setup
3.1 LLMs evaluated
We evaluate two closed models(GPT-4o-mini and GPT-4o) and four open LLMs, including Aya-101 (ustun2024ayamodelinstructionfinetuned), MetaMath 7B, MetaMath-Mistral 7B (yu2023metamath), and Gemma 2 9B (gemmateam2024gemma2improvingopen). Aya-101 is only evaluated for the AfriXNLI as it is one of the strongest open model baseline for low-resource languages.
3.2 Baselines Method Compared
We evaluate our method against five strong baselines: three model merging approaches—LangBridge (yoon-etal-2024-langbridge), MindMerger (huang2024mindmergerefficientboostingllm), LayAlign (ruan2025layalignenhancingmultilingualreasoning) and two methods leveraging in-domain question translation data, QAlign-LoRA (zhu2024questiontranslationtrainingbetter) and SLAM (fan2025slamefficientmultilingualreasoning). LangBridge injects a frozen multilingual encoder before a decoder-only LLM, replacing token embeddings with linearly projected encoder states ($\approx$3M parameters) learned from 100k English instructions, and inference cost equals that of the LLM plus the multilingual encoder. QAlign-LoRA first aligns multilingual prompts via question translation and then fine-tunes lightweight LoRA adapters (<1% parameters) on English tasks. SLAM fine-tunes the six lowest feed-forward layers ($\approx$7% parameters) identified as language-sensitive on mixed English and translated data, freezing higher layers and attention weights. MindMerger connects a frozen multilingual encoder to the LLM through a two-layer MLP head ($\approx$9M parameters), trained first on translation pairs with English references, then on translated QA pairs while keeping all base weights fixed. LayAlign integrates encoder features into multiple LLM layers via a trainable layer-wise aligner, trained in two stages using translation and task data while keeping both encoder and LLM frozen.
For the encoder component in multilingual‐integration approaches, we couple both MindMerger and Merlin with the distilled 600M‐parameter NLLB encoder (nllbteam2022languageleftbehindscaling). For LangBridge, we follow its original setup: AfriTeva-v2-Large (oladipo-etal-2023-better) is used on AfriMGSM and AfriXNLI, whereas mT5-XL is employed for MGSM, because the publicly released LangBridge implementation only supports encoders that subclass the XXXEncoderModel.
3.3 Languages and Tasks
Our evaluation spans four multilingual benchmarks. AfriMGSM (adelani2024irokobenchnewbenchmarkafrican) extends MGSM with grade-school mathematics problems translated into 16 African languages, providing fine-grained coverage of more low-resource settings. AfriXNLI (adelani2024irokobenchnewbenchmarkafrican) offers natural-language inference data in the same 16 languages by translating ten domain splits from the original XNLI. The canonical MGSM corpus (shi2022languagemodelsmultilingualchainofthought) supplies math QA instances across eight mid-resource languages. Finally, MSVAMP (chen2024breakinglanguagebarriersmultilingual) introduces multilingual variants of SVAMP with 1000 algebra-word problems in eleven languages, testing cross-lingual reasoning beyond GSM-style arithmetic.
3.4 Experimental Settings
Training Datasets
Sub-stage Ia (Mapping). For every language we sample 9 000 general-domain English–target sentence pairs from the NLLB training corpus (nllbteam2022languageleftbehindscaling). These bitext pairs supply a coarse bilingual anchor for the connector.
Sub-stage Ib (Question Alignment).
To align non-English prompts with English semantics, we translate 20 000 English questions with NLLB 200-3.3B and sample 3 000 translated pairs per language. For the mathematics track (AfriMGSM and MGSM) the source questions come from MetaMathQA (yu2024metamathbootstrapmathematicalquestions). For the NLI track (AfriXNLI) we instead sample premise–hypothesis pairs from translated MultiNLI (williams2018broadcoveragechallengecorpussentence), concatenate them as the “question”, translate, and again retain 3 000 pairs.
Sub-stage Ic (Augmentation).
For each target language we extract a distinct set of 3 000 question–answer pairs from the training split of the task corpus: translated MetaMathQA for the math benchmarks (AfriMGSM, MGSM, and MSVAMP) and translated MultiNLI for the NLI benchmark (AfriXNLI).
Task-Specialization Stage.
We fine-tune the LLM (via PEFT) on a second sample of 3 000 training instances per language drawn from the same translated sources as above. Using a separate set of examples helps reduce the risk of overlap with Sub-stage Ic and ensures there is no data leakage into the evaluation sets. [^1] The detailed training configuration is stated on Appendix
Map Align Aug Spe eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul Avg $\checkmark$ $\times$ $\checkmark$ $\times$ 58.0 50.8 45.2 32.4 44.8 34.4 43.2 38.4 36.4 40.0 44.8 36.8 53.2 24.4 21.2 33.6 34.8 40.4 39.60 $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 82.8 78.0 52.8 34.0 60.4 50.4 52.0 47.2 46.8 24.0 52.8 53.2 73.6 34.8 23.6 43.2 53.6 52.4 50.87 $\checkmark$ $\times$ $\checkmark$ $\checkmark$ 86.0 81.2 59.6 37.2 59.6 50.8 52.8 45.2 44.4 30.4 54.0 53.6 73.2 36.8 23.6 48.8 53.6 54.0 52.49 $\checkmark$ $\checkmark$ $\times$ $\checkmark$ 77.2 64.0 26.8 16.0 30.8 20.0 28.0 19.2 20.3 13.9 24.7 24.5 33.5 16.8 10.8 22.3 24.5 24.7 27.67 $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 80.0 73.2 57.2 35.6 56.8 48.0 57.2 44.4 45.2 30.4 57.6 55.2 74.0 34.0 24.0 48.4 45.2 50.4 50.93 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 82.0 74.0 59.2 41.2 60.8 52.0 52.0 49.6 46.8 32.0 58.8 56.0 76.8 37.6 26.0 51.2 54.0 55.6 53.64
Table 4: Ablation study: each row removes one stage ($\times$) while keeping the others ($\checkmark$). Map = General Bilingual Mapping; Align = Question Alignment; Aug = Task-aware Augmentation; Spe = Task Specialization Stage. The first row corresponds to the MindMerger experiment setting.
4 Experimental Results
We assessed Merlin on both math reasoning and NLI benchmarks, comparing it against open LLMs, recent multilingual integration methods, and the strongest available closed models. For math reasoning, we use two datasets: MSGM and MSVAMP. To assess performance on low-resource African languages across both tasks, we employ AfriMGSM and AfriXNLI. All results are reported as exact-match or accuracy; higher is better.
Merlin achieved SOTA on standardized math reasoning benchmarks
Table 1 presents results on MGSM and its algebraic variant MSVAMP. Using the identical Gemma-2 9B backbone, Merlin achieves $76.2\,\%$ average accuracy on MGSM, exceeding MindMerger by +2.8+2.8 points and SLAM by +22.8+22.8 points. [^2] LangBridge and QAlign scores are omitted in Table 1 because their publicly released implementations do not support the full language set. Improvements are consistent across both mid-resource languages (Bn/Th/Sw, +13.7+13.7 over Gemma-2) and high-resource languages (+7.7+7.7 over MetaMath-Mistral). On MSVAMP, Merlin retains its lead with $79.2\,\%$ average accuracy, outperforming MindMerger by +0.9+0.9 points and exceeding the best open baseline by more than $15$ points. These gains indicate that the two-stage alignment not only transfers arithmetic skills but also generalises to algebraic word problems.
Merlin achieved a wider performance gap in math reasoning for LRLs
Table 2 focuses on sixteen truly low-resource African languages. With the strong Gemma-2 9B+NLLB-600M configuration, Merlin attains $50.6\,\%$ average accuracy, surpassing MindMerger by +12.9+12.9 points, SLAM by +11.6+11.6 points, and QAlign-LoRA by +16.9+16.9 points. Notably, Merlin outruns the closed GPT-4o-mini by +15.2+15.2 points and comes within $2.0$ points of GPT-4o despite the latter’s much larger parameter budget. Ablation confirms (Table 4) that the task-aware augmentation sub-stage is decisive, yielding gains of more than ten points when added to the connector training.
Reasoning capabilities generalizes to NLU tasks
Table 3 reports accuracy on sixteen African languages plus English and French on NLI task. Merlin attains $68.7\,\%$ average accuracy on the African subset, improving on MindMerger by +1.6+1.6 points and on the base Gemma–2 9B model by +28.4+28.4 points, and marginally surpassing the closed GPT‑4o ($64.3\,\%$). The gains are most pronounced for Amharic (+5.6), Oromo (+4.6), Hausa (+2.5), and Kinyarwanda (+2.0), confirming the value of the cross‑lingual mapping curriculum in truly low‑resource settings. The overall improvement across thirteen of the sixteen African languages indicates that Merlin successfully redistributes modelling capacity toward languages that benefit the most from additional cross‑lingual supervision, while maintaining competitive performance on those already well supported.
Figure 2: Comparison of Merlin performance across five different multilingual encoders and Gemma 2 9B LLM.
5 Analysis of Language alignment and Effects on multilingualism
Figure 3: T-SNE visualizations of sentence embeddings for mT5-xl, AfriTeVa-V2-Large, and NLLB200-distilled-1.3B, showing how LRL data impacts cross-lingual alignment. More data reduces clustering, bringing low-resource languages closer to English, enhancing transfer learning.
Effectiveness in Low and Mid-resource reasoning Settings
[…]
Encoders rich in LRL data sharpen cross-Lingual alignment
A key factor in multilingual model merging is how different encoder choices affect the representation of LRLs in the mapping layer. To isolate this effect, we kept the Merlin architecture unchanged across all experiments and varied only the multilingual encoder. As shown in Figure 3, when the proportion of LRL text is small, non-English data—especially from LRLs—forms distinct clusters far from English in the LLM’s original embedding space. This separation prevents the model from sharing or transferring internal reasoning paths learned predominantly through English data. Conversely, incorporating more low-resource data enables the mapping layer to align different language embeddings, including those from LRLs, more closely with English. By bringing LRLs into the same representational space as English, the LLM can more effectively leverage its inherent reasoning capabilities for LRL tasks, rather than leaving them isolated at the margins of the embedding space. Consequently, increasing low-resource language datasets during encoder training promotes cross-lingual knowledge transfer and enhances representation consistency across languages.
Middle Decoder Layers Reach Peak Cross-Lingual Alignment
Figure 4 reports sentence-level retrieval@5 for every decoder layer of Gemma2-9B-It, MindMerger, and Merlin.
Figure 4: Layer-wise cross-lingual retrieval@55. Scores are averaged over English and 16 African languages. Merlin achieves the highest mid-layer alignment, whereas MindMerger peaks only in the final decoder blocks.
Retrieval@5 is computed by taking each English sentence in the FLORES-200 dev set as a query, ranking all candidate sentences in the paired target language by the cosine similarity of their layer-wise embeddings, and then checking whether the gold translation appears among the top-5 nearest neighbours. The score—averaged over 200 sentence pairs and 15 African languages—therefore reflects the probability that a representation keeps the correct cross-lingual match within a five-item shortlist, providing a direct proxy for how language-agnostic the layer’s embedding space is.
From layers 18–28 Merlin reaches its peak sentence–level retrieval@5 (~0.30). The rise coincides with the PEFT fine–tuning in Stage II, which adjusts a small set of low-rank weights on top of the frozen LLM while keeping the connector $f_{\sigma^{\star}}$ fixed. This adaptation strengthens the multilingual abstractions that naturally emerge in the decoder’s middle blocks—the region reported to be the most language-agnostic once parallel supervision is available (liu2025middlelayerrepresentationalignmentcrosslingual). As a result, non-English queries are already embedded in the English reasoning sub-space before the later, generation-oriented layers are reached, providing a solid cross-lingual scaffold for downstream mathematical reasoning tasks such as AfriMGSM.
6 Related Work
LLMs in LRLs
LRLs often suffer from limited annotated data and high-quality machine translation resources, creating significant challenges for NLP models (magueresse2020lowresourcelanguagesreviewpast; lee_pre-trained_2022; app13158566; deshpande2024chainoftranslationpromptingcotrnovel; qin2024multilinguallargelanguagemodel; tonja-etal-2024-nlp). Moreover, these languages can exhibit complex morphological structures and orthographic variations (ghosh2024morphologybasedinvestigationpositionalencodings; nzeyimana2024lowresourceneuralmachinetranslation; lopo2024constructingexpandinglowresourceunderrepresented; issaka2024ghanaiannlplandscapelook; lusito2023textnormalizationlowresourcelanguages), posing further challenges. While multilingual models can handle many languages, they frequently underperform in truly low-resource settings and may even trail behind strong monolingual baselines (yoon-etal-2024-langbridge; du2024chinesetinyllmpretraining; blevins-etal-2024-breaking). Existing efforts to boost LRL performance focus primarily on dataset creation (adelani-etal-2021-masakhaner; muhammad2022naijasentinigeriantwittersentiment; adelani-etal-2024-sib; yong2024lexcgengeneratingdataextremely) or adaptation of pretrained models (alabi-etal-2022-adapting; wu2024adaptinglargelanguagemodels; csaki2023efficientlyadaptingpretrainedlanguage; schmidt-etal-2024-self), yet these strategies remain constrained by data scarcity and fail to fully address complex linguistic properties (urbizu-etal-2023-scaling; wang2023challengingbenchmarklowresourcelearning).
Multilingual Reasoning.
Recent efforts to enhance multilingual reasoning in LLMs include carefully devised prompt engineering to facilitate multi-step inference or chain-of-thought reasoning (huang-etal-2023-languages; qin-etal-2023-cross), although these techniques often lose effectiveness with smaller open-source models (touvron2023llama2openfoundation). Another approach involves supervised fine-tuning on translated data (chen2024breakinglanguagebarriersmultilingual; zhu2024questiontranslationtrainingbetter), which infuses language or task-specific knowledge into the model. Building upon this, fan2025slamefficientmultilingualreasoning extend the work of zhu2024questiontranslationtrainingbetter by optimizing both the training parameters and the training stages focusing on language-specific neurons. Their improvements lead to higher average performance on multilingual mathematical reasoning benchmarks. However, reliance on machine translation can introduce errors (shi2022languagemodelsmultilingualchainofthought) and may overlook the LLM’s built-in multilingual capabilities. To circumvent these issues, MindMerger (huang2024mindmergerefficientboostingllm) sidesteps autoregressive translation by directly integrating an external multilingual encoder, preserving the LLM’s internal parameters.
7 Conclusion
This work introduced Merlin, a lightweight framework for strengthening multilingual mathematical reasoning and natural‑language inference in large language models. A two‑stage curriculum first learns a compact connector that projects multilingual encoder states into the frozen LLM embedding space, then refines the decoder through DoRA‑based task specialization. The approach preserves the original English reasoning capabilities while substantially improving performance in both mid‑ and low‑resource languages.
Our analysis shows that model merging enhances cross-lingual alignment across layers compared to using LLMs alone. In particular, our framework, Merlin, improves alignment within the model’s internal reasoning subspace, yielding stronger performance on reasoning tasks. We also find that the choice of encoder is critical for low-resource languages: encoders with greater exposure to LRL data produce representations that align more closely with the LLM’s internal language (often English), thereby supplying richer cross-lingual signals for model merging.
Empirical evaluation on four benchmarks demonstrates that Merlin outperforms strong open‑source systems—including MindMerger, SLAM, and LangBridge—on MGSM, MSVAMP, AfriMGSM and AfriXNLI, and narrows the gap to closed models such as GPT‑4o.
8 Limitations
Merlin relies on automatically translated corpora for all three training stages, assuming that these translations form a reliable semantic bridge between the target language and English. In practice, translation quality is uneven. Even with a strong multilingual model such as the distilled NLLB 600M encoder, we observe systematic errors for certain languages—most notably Oromo—where mistranslations propagate through the mapping layer and erode downstream accuracy. Because the present pipeline admits every machine-translated sentence without verification, noisy lexical choices and hallucinated numerals enter the training signal unchecked. Mitigating this weakness will require explicit data filtering.
A second limitation is the narrow scope of our evaluation. The experiments target two problem families—grade-school mathematics and natural-language inference—selected for their diagnostic value in cross-lingual reasoning. Tasks that demand other forms of structured reasoning, such as program synthesis, commonsense QA, or multi-document summarization, remain unexplored.
Finally, our methodology trains a separate Merlin instance for each benchmark. This design clarifies the contribution of the mapping curriculum but precludes parameter sharing across tasks and prevents the model from exploiting potential synergies between, for example, symbolic arithmetic and commonsense inference. A unified multi-task variant would be more practical in deployment and could yield additional gains through cross-task transfer, although its optimization dynamics are not yet understood.
Appendix A Dataset-Size Ablation
With the full sweep of dataset sizes in Table 5, a clear elbow appears at 1 000 translated sentence pairs per language and per stage. At that level Merlin attains its highest averages—78.3 % on MGSM and 80.4 % on MSVAMP—while using only one-sixth of the data required by the largest setting. A modest increase to 1 500 pairs produces small, mixed fluctuations, and every larger tier either plateaus or declines, a pattern we attribute to domain mismatch and over-fitting of the lightweight connector once the easiest bilingual correspondences have been captured. Even the 500-pair condition is competitive (75.2 % / 80.1 %), suggesting that the first few hundred high-quality translations already establish a usable cross-lingual bridge. In terms of compute, the 1 000-pair configuration is exceptionally light: all three connector stages plus DoRA adaptation complete in just under one hour on a single A100-80 GB GPU (batch size = 1), confirming that Merlin can be trained end-to-end in resource-constrained settings without sacrificing downstream accuracy.
Model / Size Bn Th Sw Ja Zh De Fr Ru Es En Mrl. Hrl. Avg
MGSM
500 72.8 73.6 71.6 72.4 73.2 74.8 75.6 79.6 76.8 82.0 72.7 76.3 75.2
1000 76.4 77.6 74.4 71.6 76.0 78.4 77.6 81.6 83.6 85.6 76.1 79.2 78.3
1500 75.6 74.4 71.6 72.8 76.4 74.4 75.6 78.8 85.2 84.8 73.9 78.3 77.0
3000 74.8 76.0 77.6 71.6 72.8 73.2 74.8 78.4 77.6 85.6 76.1 76.3 76.2
4500 74.8 77.6 72.8 74.0 74.4 75.6 75.6 78.4 82.8 84.0 75.1 77.8 77.0
9000 74.8 78.4 74.8 74.0 72.4 76.0 74.8 77.6 81.2 83.6 76.0 77.1 76.8
18000 76.0 79.6 79.2 73.2 72.8 76.8 79.2 78.0 82.8 83.6 78.3 78.1 78.1
MSVAMP
500 72.6 75.3 77.8 80.7 81.2 82.8 82.5 81.0 83.1 83.5 75.2 82.1 80.1
1000 73.0 75.7 79.4 79.7 82.2 82.9 82.7 81.6 82.8 84.4 76.0 82.3 80.4
1500 71.8 76.5 78.4 81.1 80.7 82.3 82.0 81.4 83.3 85.2 75.6 82.3 80.3
3000 72.1 76.1 77.4 78.7 79.8 81.5 81.8 79.8 81.8 83.3 75.2 81.0 79.2
4500 70.1 75.6 78.9 78.7 79.1 80.7 80.9 79.8 81.5 83.3 74.9 80.6 78.9
9000 71.6 76.0 77.1 78.7 79.5 81.3 81.3 80.4 81.6 81.9 74.9 80.7 78.9
18000 71.4 74.7 74.6 77.1 77.2 82.4 80.3 79.1 80.9 85.3 73.6 80.3 78.3
Table 5: Merlin performance on MGSM and MSVAMP across synthetic‑training set sizes. Bn/Th/Sw are mid‑resource (MRL); the remaining seven languages are high‑resource (HRL).
Appendix B Complete Encoder Ablation Study
Table 6 compares four multilingual encoders (mT5-Large, mT5-XL, AfriTeVa-v2-Large, and NLLB-600 M distilled) while keeping the Merlin decoder, connector, and all training settings fixed. Two evaluation suites are considered: AfriXNLI (classification) and AfriMGSM (mathematical reasoning).
On AfriXNLI, the distilled NLLB encoder delivers the highest average accuracy ($55.6\,\%$), improving over the best mT5 variant by more than 30 points and over AfriTeVa-v2 by 20 points. Gains are especially large for the most data-scarce languages such as Amharic (+49.5 vs. mT5-Large) and Oromo (+27.5 vs. AfriTeVa-v2), confirming that broad pre-training coverage is critical for cross-lingual alignment.
For AfriMGSM, NLLB again attains the highest average ($51.5\,\%$), edging out mT5-XL by two points. While mT5-XL excels on high-resource pivots (English, French) and on Swahili, it underperforms on several low-resource languages (e.g. Ewe, Wolof). AfriTeVa-v2 provides competitive results on Swahili and Sotho but lags elsewhere, reflecting its narrower linguistic coverage.
Overall, the results highlight that encoders trained on a broad, balanced corpus—such as NLLB—yield the most reliable cross-lingual representations when coupled with Merlin, particularly in low-resource African settings. Models with less coverage (mT5-XL) or regional focus (AfriTeVa-v2) may excel on individual languages but fall short in overall robustness.
Model eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul Avg
AfriXNLI
mT5-Large 22.8 37.8 9.7 40.0 33.5 14.5 29.0 22.0 13.5 25.7 35.3 39.0 38.2 9.5 25.7 16.8 19.0 17.0 24.3
mT5-XL 24.8 24.8 18.3 25.0 21.3 7.0 25.3 26.8 13.0 9.0 21.0 19.2 16.5 12.7 22.2 14.3 9.8 20.2 17.6
AfriTeVa V2 Large 61.0 66.0 14.3 45.2 42.3 31.5 40.3 29.3 31.2 30.7 44.5 42.8 40.7 26.2 34.0 39.5 33.0 38.3 35.2
NLLB 600M distilled 83.2 62.0 69.2 47.9 69.0 43.9 68.4 32.2 47.2 53.2 63.4 66.7 60.5 45.0 52.5 56.7 54.5 58.9 55.6
AfriMGSM
mT5-Large 80.0 68.0 49.6 13.6 54.0 27.6 40.0 23.2 31.2 18.8 40.0 49.2 64.4 12.4 7.6 39.6 27.6 35.6 34.6
mT5-XL 82.8 80.0 62.0 28.4 64.0 50.4 52.4 41.2 41.2 42.8 62.0 59.2 79.2 24.0 15.2 49.2 47.2 55.2 49.5
AfriTeVa V2 Large 85.2 77.2 60.0 11.6 62.0 47.6 52.4 31.6 40.8 51.2 53.6 55.2 77.2 11.6 8.8 46.8 53.6 54.0 46.0
NLLB 600M distilled 82.0 74.0 59.2 41.2 60.8 52.0 52.0 49.6 46.8 32.0 58.8 56.0 76.8 37.6 26.0 51.2 54.0 55.6 51.5
Table 6: Accuracy of four multilingual models on AfriXNLI and AfriMGSM. Avg is computed over the 16 African languages (all columns except eng and fra).
Appendix C Ablation of Training Mapping Layers
Table 7 varies the expressiveness of the mapping head while holding every other component of Merlin fixed. A linear projection or a single hidden layer yields only modest averages and large per-language variance. Adding a second hidden layer (9.45 M parameters) stabilises performance, but the decisive improvement comes from inserting a residual skip: the resulting two-layer Residual MLP attains 77.6 % on MGSM and 81.1 % on MSVAMP, the best figures in both suites. Expanding to three hidden layers (13.6 M parameters) reverses the gain, dropping accuracy by 1–2 points and confirming that extra depth overfits the limited translation supervision.
The residual two-layer MLP therefore offers the strongest accuracy–capacity trade-off, adding fewer than ten million parameters while outperforming shallower and deeper variants across mid- and high-resource languages. At the same time the gap between 77.6 % and the theoretical ceiling on MGSM suggests room for architectural exploration—e.g. gated skips, attention-based adapters, or language-conditioned hyper-networks—that could push the mapping layer beyond its current residual design.
Mapping head # Parm Bn Th Sw Ja Zh De Fr Ru Es En Mrl. Hrl. Avg
MGSM
Linear 3.68 M 72.4 75.2 70.4 75.2 74.0 79.2 77.2 81.2 80.8 85.2 72.7 78.9 77.08
1-Layer MLP 3.68 M 65.6 68.8 71.2 68.8 62.8 71.6 73.6 78.0 77.2 82.0 68.5 73.4 71.5
2-Layer MLP 9.45 M 74.8 76.0 77.6 71.6 72.8 73.2 74.8 78.4 77.6 85.6 76.1 76.3 76.2
3-Layer MLP 13.64 M 74.8 76.4 69.6 70.0 71.2 74.0 72.0 76.8 79.2 79.6 73.6 74.7 74.4
Residual MLP 9.45 M 75.2 80.0 77.2 73.6 70.8 78.8 78.0 79.2 82.4 80.8 77.5 77.7 77.6
MSVAMP
Linear 3.68 M 72.8 75.2 77.8 79.5 79.9 83.7 81.4 80.3 82.0 83.0 75.3 81.4 79.6
1-Layer MLP 3.68 M 69.4 75.8 77.3 79.3 79.9 81.1 81.3 80.0 81.7 82.4 74.2 80.8 78.8
2-Layer MLP 9.45 M 72.1 76.1 77.4 78.7 79.8 81.5 81.8 79.8 81.8 83.3 75.2 81.0 79.2
3-Layer MLP 13.64 M 71.0 74.0 76.7 77.3 76.1 80.8 82.0 78.7 80.6 81.5 73.9 79.6 77.9
Residual MLP 9.45 M 72.7 77.6 80.7 79.6 81.0 84.8 83.3 83.2 83.7 84.5 77.0 82.9 81.1
Table 7: Effect of mapping-head capacity on MGSM and MSVAMP. Mrl. = mean over mid-resource languages (Bn, Th, Sw); Hrl. = mean over the high-resource set; “–” indicates the score is not available.
Appendix D Ablation of Training Adapters
Table 8 contrasts two parameter-efficient adapters applied to the decoder. LoRA (hu2022lora) injects trainable low-rank update matrices into the linear weights, while all original parameters stay frozen. DoRA (liu2024doraweightdecomposedlowrankadaptation) keeps the same low-rank structure but factorises each weight into a direction and a magnitude, yielding the same footprint yet a more explicit decomposition.
With rank and parameter count held constant, the two adapters behave nearly identically: average accuracy on MGSM differs by only 0.2 pp, and the MSVAMP results coincide after rounding. Performance across mid- and high-resource language subsets mirrors this pattern, indicating that adapter choice has minimal impact under our training recipe.
Adaptation # Parm Bn Th Sw Ja Zh De Fr Ru Es En Mrl. Hrl. Avg
MGSM
LoRA 18.4 M 75.2 76.4 77.2 71.6 72.4 72.4 75.6 79.2 78.0 85.6 76.3 76.4 76.4
DoRA 18.4 M 74.8 76.0 77.6 71.6 72.8 73.2 74.8 78.4 77.6 85.6 76.1 76.3 76.2
MSVAMP
LoRA 18.4 M 71.6 76.2 77.3 78.0 79.1 81.6 82.1 80.1 82.2 83.6 75.0 81.0 79.2
DoRA 18.4 M 72.1 76.1 77.4 78.7 79.8 81.5 81.8 79.8 81.8 83.3 75.2 81.0 79.2
Table 8: Ablation of LoRA vs. DoRA adapters for Merlin (Gemma-2 9 B + NLLB-600 M). Mrl. = mean over mid-resource languages (Bn, Th, Sw); Hrl. = mean over seven high-resource languages; averages are arithmetic means of the shown scores.
Appendix E AmericasNLI Evaluation
To assess whether the mapping curriculum generalises beyond the African and European test beds, we evaluate Merlin on AmericasNLI, a three-way NLI benchmark that targets Aymara, Guarani, and Quechua. For Merlin and the MindMerger baseline, we follow the same data pipeline as for AfriXNLI: 3,000 English premise–hypothesis pairs are sampled from MultiNLI, translated into each target language with NLLB 600m distilled, and used as supervision for Stage I.b, Stage I.c, and Stage II. [^3] No task-specific data from the original AmericasNLI development set is seen during training. The non-adapted LLaMAX2-7B-XNLI (lu-etal-2024-llamax) serves as the baseline decoder-only model. Results are summarised in Table 9.
Model Aymara Guarani Quechua Avg.
LLaMAX2 XNLI (base) 41.6 43.6 40.7 42.0
MindMerger 48.8 48.7 48.3 48.6
Merlin 48.4 49.3 50.5 49.4
Table 9: Accuracy (%) on AmericasNLI. Average is computed over the three indigenous languages.
Merlin outperforms the base model by +7.4+7.4 accuracy points on average and still outperform past MindMerger by +0.8+0.8 points. The consistent improvement across all three languages suggests that the cross-lingual alignment learned from African and Indo-European data transfers to typologically distant language families in the Americas.
Appendix F Training Configurations
All parameter-efficient runs (Merlin, QAlign-LoRA) employ rank-16 LoRA/DoRA adapters injected into the self‐attention q_proj and v_proj projections:
Hyper-parameter Value
Rank $r$ 16
LoRA / DoRA $\alpha$ 32
Dropout 0.05
Bias none
Target modules q_proj, v_proj
QAlign and SLAM adopt identical settings from the released code  [^4] https://github.com/NJUNLP/QAlign,https://github.com/fmm170/SLAM. For LangBridge we follow the original implementation, using a single linear projection layer and sampling 100k examples from MetaMathQA training set
Mathematical Reasoning (MGSM, MSVAMP, AfriMGSM)
Gemma-2 backbone <bos><start_of_turn>user Below is an instruction that describes a task. Write a response that appropriately completes the request. {query} Let’s think step by step. <end_of_turn><start_of_turn>model
MetaMath-style backbones Below is an instruction that describes a task. Write a response that appropriately completes the request. ### Instruction: {query} ### Response: Let’s think step by step.
Natural-Language Inference (AfriXNLI)
All backbones Premise: {sentence1} Hypothesis: {sentence2} Label:
Table 10: Prompt templates used for Merlin, MindMerger and QAlign. Curly braces denote insertion points for the instance text.
Both Merlin and MindMerger are trained on identical corpora. Stage I.a uses 9,000 generic English–target sentence pairs per language drawn from the NLLB-200 corpus. Stage I.b is supplied with up to 3,000 automatically translated questions per language, also generated with NLLB. For Stage I.c and the subsequent Task-Specialisation stage, task-specific data are employed: the MGSM track relies on the translated MetaMathQA split, whereas the AfriXNLI track is trained with translated MultiNLI.
Appendix G Prompt Templates
LangBridge and SLAM are executed through lm-eval-harness; their default prompts are therefore retained and not repeated here. Table 10 shows the exact text used by Merlin, MindMerger, and QAlign for the two evaluation tracks.